Let ?(?, ?) = (? − ?)? + 3? ? + ?^2?. Find an equation of the tangent plane to the surface of ?(?, ?) at the point (−1, 2, 1).
Answer #1

8(4,0) = (u-v) it 3 vj+ u²k ru a alu-u)i + a 1 (31) j + d (14²) ru = irojtzuk 8v = a (u-u)i + 2(30)j + d(u²) k y = -1 +3j tok